# AutoDroid-V2: Boosting SLM-based GUI Agents via Code Generation

### v2

@@ Abstract.
Large language models (LLMs) have brought exciting new advances to mobile UI agents, a long-standing research field that aims to complete arbitrary natural language tasks through mobile UI interactions. However, existing UI agents usually demand high reasoning capabilities of powerful large models that are difficult to be deployed locally on end-users’ devices, which raises huge concerns about user privacy and centralized serving cost. One way to reduce the required model size is to customize a smaller domain-specific model with high-quality training data, e.g. large-scale human demonstrations of diverse types of apps and tasks, while such datasets are extremely difficult to obtain. Inspired by the remarkable coding abilities of recent small language models (SLMs), we propose to convert the UI task automation problem to a code generation problem, which can be effectively solved by an on-device SLM and efficiently executed with an on-device code interpreter. Unlike normal coding tasks that can be extensively pretrained with public datasets, generating UI automation code is challenging due to the diversity, complexity, and variability of target apps. Therefore, we adopt a document-centered approach that automatically builds fine-grained API documentation for each app and generates diverse task samples based on this documentation. By guiding the agent with the synthetic documents and task samples, it learns to generate precise and efficient scripts to complete unseen tasks. Based on detailed comparisons with state-of-the-art mobile UI agents, our approach effectively improves the mobile task automation with significantly higher success rates and lower latency/token consumption. Code will be open-sourced.
@@
copyright: acmcopyrightjournalyear: 2023copyright: rightsretainedconference: Preprint; price: ;submissionid: #128
@@ 1. Introduction
Automatically controlling mobile devices through natural language has long been a goal for mobile developers and researchers (Bolt, 1980; Azim et al., 2016; Li et al., 2024). Agents powered by Large Language Models (LLMs) and Vision Language Models (VLMs, both LLM and VLM are refered as LLM for simplicity in this paper) have emerged as promising solutions for automating flexible and complex user tasks (Yuan et al., 2024; Wen et al., 2024; You et al., 2024; Wang et al., 2024c; Lee et al., 2023; Liu et al., 2024; Zhang et al., 2024a; Wen et al., 2023). These agents function by interpreting user instructions, analyzing the graphical user interface (GUI) states, and predicting UI actions (e.g., touch, scroll, input text, etc.) to be performed in the target apps iteratively. Most of these LLM-based GUI agents operate in a step-by-step manner, i.e. querying the LLM at each GUI state for dynamic decision and reflection. They typically include the user’s task and the observed GUI state (e.g., a screenshot or a text-based view hierarchy) in the LLM prompt, requesting the model to generate the required UI action. This iterative process continues until the task is completed. For this reason, we refer to these conventional designs as “Step-wise GUI agents” in this paper.
@@
While step-wise GUI agents hold great potential, they mostly rely on powerful reasoning and reflection abilities, which are dominated by large cloud-based models (e.g. GPT-4o, Claude, etc.). Due to the high computational cost and/or closed-source nature of these models, they can hardly be deployed to mobile devices. Such a situation can lead to significant privacy and security risks of sharing each personal GUI page to the cloud (Zhang et al., 2024c; Ma et al., 2024; Zhang et al., 2024d), high user-side traffic consumption, and high server-side centralized serving costs, hindering large-scale deployment of GUI agents.
@@
Among the possible countermeasures to this problem, small language models (SLMs) (There is no clear boundary between LLM and SLM. In this paper we use SLM to represent the LLMs that can be deployed on the device.) and efficient inference systems have emerged as promising directions to enable on-device deployment of more powerful models (Van Nguyen et al., 2024; Xu et al., 2024; Lu et al., 2024). The current capabilities of these SLMs are still far from their cloud-based counterparts, and the capability gap is anticipated to keep existing in the near future due to the scaling laws (Kaplan et al., 2020). Customizing a domain-specific SLM can also be a potential solution. However, training a domain-specific model requires a large-scale, high-quality labeled dataset covering diverse usage scenarios, which are difficult to obtain in the UI task automation domain. Most existing mobile task automation datasets (Burns et al., 2022; Zhang et al., 2024b; Rawles et al., 2024; Sun et al., 2022) only have hundreds of human demonstrations. The largest dataset AitW (Rawles et al., 2023) has 715k crowd-sourced task demonstrations, while the quality is low for most demonstrations (We compare the agents trained with such datasets in Section 4.2). Meanwhile, the LLM abilities used in existing GUI agents are diverse, which include not only simple GUI grounding abilities that can be offered by a small tailored model  (Cheng et al., 2024; Hong et al., 2024; Gou et al., 2024), but also general reasoning and reflection abilities that can hardly be distilled to a much smaller model (Zhang et al., 2024a; Wang et al., 2024b; Bishop et al., 2024).
@@
To this end, we introduce AutoDroid-V2 to investigate how to build a powerful GUI agent upon the coding capabilities of SLMs. The key difference of AutoDroid-V2 is to complete GUI tasks by generating and executing multi-step scripts based on user instructions, rather than deciding one action at a time. We refer to this approach as the “Script-based GUI Agent” to distinguish from traditional step-wise GUI agents, as illustrated in Figure 1. Such a script-based agent can effectively address the aforementioned two problems of step-wise approaches: (1) Efficiency: Agents can generate a single script for a series of GUI actions to complete a task based on the user task, significantly reducing query frequency and consumption. (2) Capability: Script-based GUI agents rely primarily on the coding ability of SLMs, which have been proven effective by numerous existing studies on lightweight coding assistants.
@@
    We investigate a new paradigm of an on-device GUI agent that programmatically interacts with GUIs, presenting a promising solution to UI task automation based on the remarkable coding capabilities of SLMs and efficient code-based task execution.
@@ 2.1. GUI-based Mobile Task Automation
The goal of mobile task automation is to automatically complete user tasks by interacting with mobile devices. Compared to previous template-based methods (e.g. Siri, Google Assistant, Cortana, etc.), where predefined templates are used to process and respond to user inputs, GUI-based mobile task automation can achieve more complex and flexible tasks because it doesn’t depend on third-party APIs or non-trivial programming efforts from developers.
@@
The input of a GUI-based task automation agent is an arbitrary task described in natural language related to a mobile app. The output is a sequence of GUI actions that can be executed on a smartphone. A task is a multi-step functionality request from the user intended for completion on a smartphone, often lacking explicit instructions. A GUI State represents the current status or condition of the user interface in a mobile app, often including an arrangement of controls depicted through images and text, typically organized as a GUI tree. A GUI element is a control in the GUI state that human users can interact with, such as a button, text box, input field, or slider. A GUI action, performable by the user or an agent on the device’s screen, is defined by a tuple (GUI element, action type, value). Action type represents how the target GUI element is manipulated. We consider four main types of smartphone interactions, including “click”, “long click”, “input”, and “swipe”.
Before the emergence of LLMs, researchers developed supervised learning techniques (Li et al., 2020; Sun et al., 2022; Burns et al., 2022) and reinforcement learning methods (Toyama et al., 2021; Humphreys et al., 2022; Li and Riva, 2021) for GUI-based mobile task automation, which struggle with the flexibility and adaptability needed for dynamic, real-world applications. The advent of LLMs brings a transformative shift in GUI-based mobile task automation (Zhang et al., 2024a; Wen et al., 2024; Lee et al., 2024; Zhang et al., 2023). Despite this progress, LLM-based GUI agents often rely on sensitive user data, such as screenshots and user interaction logs, which may need to be transmitted to remote servers for processing. This raises significant privacy concerns, such as potential data breaches and misuse of personal information (Zhang et al., 2024a). Moreover, the high user-side traffic and server-side costs present barriers to large-scale deployment. Therefore, it is important to introduce on-device LLM-powered GUI agents for mobile task automation.
2.2. Current Practice and Limitations
Table 1. Average latency (per task) and success rate of GUI agents based on LLM (GPT-4o) and SLM (Llama3.1-8B) on DroidTask dataset.
Model Metric Mind2Web AutoDroid
GPT-4o Success Rate 65.2% 62.7%
Latency (s) 26.5 24.4
Llama3.1-8B Success Rate 34.4% 43.9%
Latency (s) 1149.1 669.2
@@ 2.2. Current Practice and Limitations
Researchers have explored various techniques to customize domain-specific SLMs for GUI-based mobile task automation. One area of focus involves enhancing the grounding capability of SLMs, which refers to locating GUI elements based on user instructions within a single GUI screen (Cheng et al., 2024; Hong et al., 2024; Gou et al., 2024). These methods gather extensive GUI-domain training data and use it to fine-tune a vision language model, thereby improving their ability to follow GUI domain-specific instructions. Other studies aim to acquire domain-specific knowledge about particular apps through exploration and use it to enhance SLMs’ task-solving capabilities (Wen et al., 2024; Bai et al., 2024).
As shown in Table 1, on-device GUI agents still exhibit lower accuracy and longer latencies compared to agents based on more powerful LLMs in the cloud. The high latency arises because model inference is computationally intensive, leading to delays on mobile devices with less computational resources, especially for prompts with long contexts. GUI agents must query the model at each step to process the GUI state and generate corresponding actions. Each GUI state typically consumes a substantial number of tokens (over 500 on average), causing significant processing delays—even for on-device SLMs. A typical task requires between 5 and 20 steps, further increasing latency.
The root cause of the performance degradation is the weak reasoning ability of SLMs. We conducted an in-depth analysis and identified three detailed reasons: (1) Limited Domain-specific Knowledge of Apps: On-device SLM-based agents often struggle to accurately interpret the specific functionality of UI components upon first encounter. For example, given the user task: Modify a contact’s phone number, agents often directly tap the phone number from the contact information page and don’t realize it needs to first navigate to the edit page. (2) Insufficient Tracking and Use of Past Actions: Agents frequently fail to maintain a coherent understanding of how previous actions contribute to the overall task. This leads to redundancy or incorrect sequences. For instance, an agent might repeatedly attempt to change the theme color or modify a contact number, not realizing that these actions have already been performed. (3) Determining the termination condition of one task: SLMs often exhibit hallucinations regarding the indicators of task completion, leading to premature termination or, conversely, endless loops of UI actions.
@@ 2.3. Script-based GUI Agents: Opportunities and Challenges
The limitations of current methods highlight the need for a more accurate and efficient approach to automating mobile UI tasks beyond traditional step-wise UI agents. Recent studies have shown that SLMs can perform well on challenging coding tasks (Zhu et al., 2024; Roziere et al., 2023), and synthesizing executable code are more advantageous than using Text/JSON actions in agents (Wang et al., 2024a). This motivates the development of a script-based on-device GUI agent to address the identified limitations, offering the following benefits:
@@ 3.1. Automated App Document Generation
The goal of automated app document generation is to capture and analyze essential GUI elements in an app, guiding the backbone language model to be aware of possible interaction use cases even without real-time observation. The input consists of random exploration traces, modeled as sequences of ¡GUI State, GUI Action¿ pairs. These sequences are often aimless and redundant, containing many repeated actions and GUI elements. The output is a concise app document containing typical GUI states, important elements, and GUI transition relationships, as shown in Figure 3. Two primary challenges arise in this process: reducing the GUI element space and modeling transition relationships between GUI elements.
@@ 3.1.1. Functionality-aware GUI States Grouping
However, layout similarity alone is insufficient. Many different GUI states in an app may share the same layout due to common development patterns. For instance, a date selection screen and a background theme selection screen in a calendar app might both use a similar layout (e.g., a popup with a series of checkboxes). Despite having the same layout, these two screens serve different functions. To address this, we introduce functionality similarity, which considers the tasks or actions that users can perform within a particular GUI state. This ensures that even if the layouts are similar, the underlying functionality is accurately captured and differentiated. AutoDroid-V2 uses GPT-4o to assess and classify each GUI group based on its functionality after the layout-based grouping is performed.
Then, we can represent each abstract GUI state $S^{(i)}$ after grouping as $(state\_name^{(i)},layout^{(i)})$, where $state\_name^{(i)}$ refers to the task or function associated with the GUI state (e.g., ”date selection” or ”theme selection”) and $layout^{(i)}$ represents the shared visual structure.
@@ 3.1.2. Abstract GUI Elements
As shown in Table 2, an abstract element comprises four necessary components: Name, Description, Identifier, Dependency along with two optional components: Action Effect, Options. For the element name, the GUI state name is used as a prefix because there may be several elements with the same name that belong to different states and function differently. For example, tapping the ’more options’ button in different GUIs can lead to different GUI states. Identifier is used during runtime to locate and interact with the desired GUI element based on the rules defined in the script. AutoDroid-V2 uses XPath (XML Path Language) as the identifier for each GUI element. Due to the frequent changes in the number and properties of GUI elements, using a single XPath to locate the GUI element can be unreliable at runtime. Therefore, we generate multiple XPaths for each GUI element, including the less changeable properties such as ‘resource_id’, and the relative position of a GUI element from the root of the GUI tree. A GUI element is matched if any one of these XPaths matches the element during runtime.
To identify abstract GUI elements through an exploration history, we enumerate all GUI elements within every GUI state and use LLMs to determine the element: Name, Description, and Options for important elements. The Effect of each element is determined by the subsequent GUI state triggered by interaction. Additionally, LLMs are used to categorize elements as dynamic or static and encapsulate them for further processing.
@@ 3.2.1. Large-scale Task & Solution Generation
20 "plan": "<a high level plan to complete the task>",
@@
Listing 1: Structure of the code generation prompt.
@@
Listing 2: LLM generated Python code for the user task.

### v3

@@ Abstract.
Large language models (LLMs) have brought exciting new advances to mobile UI agents, a long-standing research field that aims to complete arbitrary natural language tasks through mobile UI interactions. However, existing UI agents usually demand powerful large language models that are difficult to be deployed locally on end-users’ devices, raising huge concerns about user privacy and centralized serving cost. Inspired by the remarkable coding abilities of recent small language models (SLMs), we propose to convert the UI task automation problem to a code generation problem, which can be effectively solved by an on-device SLM and efficiently executed with an on-device code interpreter. Unlike normal coding tasks that can be extensively pre-trained with public datasets, generating UI automation code is challenging due to the diversity, complexity, and variability of target apps. Therefore, we adopt a document-centered approach that automatically builds fine-grained API documentation for each app and generates diverse task samples based on this documentation. By guiding the agent with the synthetic documents and task samples, it learns to generate precise and efficient scripts to complete unseen tasks. Based on detailed comparisons with state-of-the-art mobile UI agents, our approach effectively improves the mobile task automation with significantly higher success rates and lower latency/token consumption. Code is open-sourced at https://github.com/MobileLLM/AutoDroid-V2.
Mobile UI Agent, Small Language Models, Mobile Task Automation, Script Generation
@@
submissionid: #128journalyear: 2025conference: The 23rd Annual International Conference on Mobile Systems, Applications and Services; June 23–27, 2025; Anaheim, CA, USAbooktitle: The 23rd Annual International Conference on Mobile Systems, Applications and Services (MobiSys ’25), June 23–27, 2025, Anaheim, CA, USAdoi: 10.1145/3711875.3729134isbn: 979-8-4007-1453-5/25/06ccs: Human-centered computing Ubiquitous and mobile computingccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Automating the control of mobile devices through natural language has long been a goal for mobile developers and researchers (Bolt, 1980; Azim et al., 2016; Li et al., 2024). Agents powered by Large Language Models (LLM) and Vision Language Models (VLM, both LLM and VLM are refered as LLM for simplicity in this paper) have emerged as promising solutions for automating flexible and complex user tasks (Yuan et al., 2024; Wen et al., 2024; You et al., 2024; Wang et al., 2024c; Lee et al., 2023; Liu et al., 2024; Zhang et al., 2024a). These agents function by interpreting user instructions, analyzing the graphical user interface (GUI) states, and predicting UI actions (e.g., touch, scroll, input text, etc.) to be performed in the target apps iteratively. Most of these LLM-based GUI agents operate in a step-by-step manner, i.e. querying the LLM at each GUI state for dynamic decision and reflection. They typically include the user’s task and the observed GUI state (e.g., a screenshot or a text-based view hierarchy) in the LLM prompt, requesting the model to generate the required UI action. This iterative process continues until the task is completed. For this reason, we refer to these conventional designs as “Step-wise GUI agents” in this paper.
@@
While step-wise GUI agents hold great potential, they mostly rely on powerful reasoning and reflection abilities, which are dominated by large cloud-based models (e.g. GPT-4o, Claude, etc.). Due to the high computational cost and/or closed-source nature of these models, they can hardly be deployed to mobile devices. This situation can lead to significant privacy and security risks of sharing each personal GUI page to the cloud (Zhang et al., 2024d; Ma et al., 2024; Zhang et al., 2024e), high user-side traffic consumption, and high server-side centralized serving costs, hindering large-scale deployment of GUI agents.
@@
Among the possible countermeasures to this problem, small language models (SLMs) (There is no clear boundary between LLM and SLM. In this paper, we use SLM to represent the LLMs that can be deployed on the device.) and efficient inference systems have emerged as promising directions to enable on-device deployment of more powerful models (Van Nguyen et al., 2024; Xu et al., 2024; Lu et al., 2024). The current capabilities of these SLMs are still far from their cloud-based counterparts, and the capability gap is anticipated to remain in the near future due to the scaling laws (Kaplan et al., 2020). Customizing a domain-specific SLM can also be a potential solution. However, training a domain-specific model requires a large-scale, high-quality labeled dataset covering diverse usage scenarios, which are difficult to obtain in the UI task automation domain. Most existing mobile task automation datasets (Burns et al., 2022; Zhang et al., 2024c; Rawles et al., 2024; Sun et al., 2022) only have hundreds of human demonstrations. The largest dataset AitW (Rawles et al., 2023) has 715k crowd-sourced task demonstrations, while the quality is low for most demonstrations (we compare the agents trained with such datasets in Section 4.2). Meanwhile, the LLM abilities used in existing GUI agents are diverse, including not only simple GUI grounding abilities that can be offered by a small tailored model  (Cheng et al., 2024; Hong et al., 2024; Gou et al., 2024), but also general reasoning and reflection abilities that can hardly be distilled to a much smaller model (Zhang et al., 2024a; Wang et al., 2024b; Bishop et al., 2024).
@@
To this end, we introduce AutoDroid-V2 to investigate how to build a powerful GUI agent upon the coding capabilities of SLMs. The key difference of AutoDroid-V2 is to complete GUI tasks by generating and executing multi-step scripts based on user instructions, rather than deciding one action at a time. We refer to this approach as the “Script-based GUI Agent”, which largely differs from traditional step-wise GUI agents, as illustrated in Figure 1. Such a script-based agent can effectively address the aforementioned two problems of step-wise approaches: (1) Efficiency: Agents can generate a single script for a series of GUI actions to complete a task based on the user task, significantly reducing query frequency and consumption. (2) Capability: Script-based GUI agents rely primarily on the coding ability of SLMs, which have been proven effective by numerous existing studies on lightweight coding assistants.
@@
    We introduce the first GUI agent driven by document-guided multi-step program generation, presenting a promising solution combining the remarkable coding capabilities of SLMs and efficient code-based task execution.
@@ 2.1. GUI-based Mobile Task Automation
The goal of mobile task automation is to automatically complete user tasks by interacting with mobile devices. A GUI-based task automation agent takes a natural language task related to a mobile app and outputs a sequence of GUI actions executable on a smartphone. A task is a multi-step functionality request from the user intended for completion on a smartphone, often lacking explicit instructions. A GUI State reflects the current status or condition of the user interface in the app, usually structured as a GUI tree of controls represented by images and text. A GUI element is a control in the GUI state that human users can interact with, such as a button, a text box, an input field, and a slider. A GUI action, performable by the user or an agent on the device screen, is defined by (GUI element, action). Action represents how the target GUI element is manipulated, including “click”, “long click”, “input ¡text¿”, and “swipe ¡direction¿”.
2.2. Related Works
LLM-based GUI Agents. With the rise of large foundation models, there has been a trend leveraging LLMs (Zhang et al., 2024b; Wen et al., 2023, 2024; Lee et al., 2024; Zhang et al., 2023b) and Vision Language Models (VLMs) (Zhang and Zhang, 2024; Yan et al., 2023; Zhang et al., 2023a) to guide mobile task automation. These models enable more flexible and scalable task automation, typically executing tasks in a step-wise manner. Some research efforts focus on code generation for web page task automation (Gur et al., 2024; Tao et al., 2024), relying primarily on the HTML to generate actions within the current screen. In contrast, AutoDroid-V2 adopts a full task-level script generation approach for mobile GUI agents. LLM-based GUI agents often rely on sensitive data—like screenshots and interaction logs—that may be sent to remote servers. This creates serious privacy risks, including data breaches and misuse (Zhang et al., 2024a). High data traffic and server costs also hinder large-scale deployment. On-device LLM-powered agents are therefore essential for secure and affordable mobile task automation.
@@ 2.2. Related Works
SLM-based GUI Agents. Researchers have explored various techniques to customize domain-specific SLMs for GUI-based mobile task automation. One area of focus involves enhancing the grounding capability of SLMs, which refers to locating GUI elements based on user instructions within a single GUI screen (Cheng et al., 2024; Hong et al., 2024; Gou et al., 2024). These methods gather extensive GUI-domain training data and use it to fine-tune a vision language model, thereby improving their ability to follow GUI domain-specific instructions. Other studies aim to acquire domain-specific knowledge about particular apps through exploration and use it to enhance SLMs’ task-solving capabilities (Wen et al., 2024; Bai et al., 2024). However, the limited reasoning abilities of SLMs and large computational resource consumption hinder their widespread use, and we try to improve the efficiency and effectiveness of these methods by introducing document-guided script generation.
@@ 2.3. Script-based GUI Agents: Opportunities and Challenges
The limitations of current methods highlight the need for a more accurate and efficient approach to automating mobile UI tasks beyond traditional step-wise UI agents. Recent studies have shown that SLMs can perform well on challenging coding tasks (Zhu et al., 2024; Roziere et al., 2023), and synthesizing executable code is more advantageous than using Text/JSON actions in agents (Wang et al., 2024a). This motivates the development of a script-based on-device GUI agent to address the identified limitations, offering the following benefits:
@@ 3.1. Automated App Document Generation
The goal of automated app document generation is to capture and analyze essential GUI elements in an app, guiding the backbone SLM to be aware of possible interaction use cases even without real-time observation. The input consists of random exploration traces, modeled as sequences of ¡GUI State, GUI Action¿ pairs. Following existing works (Wen et al., 2024; Zhang et al., 2023a; Lee et al., 2024), we suppose that the exploration traces can be generated by GUI testing tools (Li et al., 2017) that explore apps randomly or through stochastic greedy depth-first strategies, or by a user trace monitor that records app usage in the background (Google, 2023). The exploration phase operates independently of user instructions, and the resulting traces are often redundant or aimless, containing many repeated actions and GUI elements. The output is a concise app document containing typical GUI states, important elements, and GUI transition relationships, as shown in Figure 3. Two primary challenges arise in this process: reducing the GUI element space and modeling transition relationships between GUI elements.
@@ 3.1.1. Functionality-aware GUI States Grouping
However, layout similarity alone is insufficient. Many different GUI states in an app may share the same layout due to common development patterns. For instance, a date selection screen and a background theme selection screen in a calendar app might both use a similar layout (e.g., a popup with a series of checkboxes). Despite having the same layout, these two screens serve different functions. To address this, we introduce functionality similarity, which considers the tasks or actions that users can perform within a particular GUI state. This ensures that even if the layouts are similar, the underlying functionality is accurately captured and differentiated. AutoDroid-V2 uses LLMs (e.g. GPT-4o) to assess and classify each GUI group based on its functionality after the layout-based grouping is performed. Specifically, for each step in the exploration history, AutoDroid-V2 constructs a prompt that includes the functionalities of previous GUI states (initially an empty list), and the current GUI state. It sends the prompt to the LLM to determine whether the current GUI state corresponds to any of the previously identified functionalities. If a match is found, the LLM maps the current GUI state to that existing functionality. If not, it generates a new name and description, which are then added to the functionality list.
Then, we can represent each abstract GUI state $S^{(i)}$ after grouping as $(functionality^{(i)},layout^{(i)})$, where $functionality^{(i)}$ refers to the task or function associated with the GUI state (e.g., ”date selection” or ”theme selection”) and $layout^{(i)}$ represents the shared visual structure.
@@ 3.1.2. Abstract GUI Elements
[…]
To identify abstract GUI elements through an exploration history, we enumerate all GUI elements within every GUI state and use LLMs to determine the element: Name, Description, and Options for important elements. The Effect of each element is determined by the subsequent GUI state triggered by interaction. Additionally, LLMs are used to categorize elements as dynamic or static and encapsulate them for further processing. Specifically, for each GUI state from one abstract GUI state, AutoDroid-V2 constructs a prompt (shortened for clarity) to query LLMs (e.g. GPT-4o) to classify the GUI elements and their properties.
1Imagine you are a UI annotator tasked with describing interactive elements of a mobile application in the form of a UI document. You are given:
2<Current UI State>
3<UI Action Performed on Current UI State> (including the functionality of the next GUI state)
4<Previously Classified UI Elements in the Same Abstract GUI State>
5<An example>
6Now, you should output all new UI elements (i.e., those not present in the previously classified elements) that can be interacted with in the current UI state.
Listing 1: Structure of GUI element classification prompt.
@@ 3.2.1. Large-scale Task & Solution Generation
20 "plan": "<a high-level plan to complete the task>",
@@
Listing 2: Structure of the code generation prompt.
@@
Listing 3: LLM generated Python code for the user task.